OpenStudy (anonymous):

I need help with a implicit diff. problem... Ln(15x^7)-3y^3=x^3e^y... i need to find y prime,.. got stuck.. any help???

OpenStudy (anonymous):

is it\[x^{3e^{y}}\]or\[x^3e^y\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

\[7/x-9y^2(dy/dx)=3x^2e^y+x^3e^y(dy/dx)\] combine the dy/dx \[(7-3x^2e^y)/x(9y^2+x^2e^y)=dy/dx\]